What will be the final temperature if 2000g of copper at 95C loses 10 kcal of heat?
Answer #1

specific heat capacity of copper, C = 0.385 J/g.oC

q = 10 kcal = 10*4.184 KJ = 41.84 KJ = 41840 J

m = 2000 g

Ti = 25.4 oC

use:

Q = m*C*(Tf-Ti)

41840 = 2000*0.385*(Tf-95)

Tf -95 = 54.34 oC

Tf = 149 oC

Answer: 149 oC